Projected Spontaneous Liabilities Smiley Corporation's current sales and partial balance sheet are shown below.

This year

Sales $ 10,000

Balance Sheet: Liabilities

Accounts payable $ 2,000

Notes payable $ 1,500

Accruals $ 1,800

Total current liabilities $ 5,300

Long-term bonds $ 2,000

Total liabilities $ 7,300

Common stock $ 1,500

Retained earnings $ 2,000

Total common equity $ 3,500

Total liabilities & equity $ 10,800

Sales are expected to grow by 14% next year. Assuming no change in operations from this year to next year, what are the projected spontaneous liabilities?

Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

Spontaneous liabilities are accounts payable and accruals

Hence projected accounts payable are

Accounts payable = 2000*114%= 2280

Accruals = 1800*114%= 2052

Total Spontaneous liabilities= $4332